# Understanding R1-Zero-Like Training: A Critical Perspective

### v1

@@ 1 Introduction
Our analysis on base models and RL suggests a minimalist recipe for R1-Zero-like training: we RL-tune Qwen2.5-Math-7B using the (unbiased) Dr. GRPO algorithm on MATH (Hendrycks et al., 2021) level 3-5 questions with the Qwen-Math template, and achieve state-of-the-art performance (Fig. 2) with only $27$ hours compute on 8×8$\times$8 × A100 GPUs. We hope our findings presented in this paper, models released, and the codebase open-sourced could benefit future research in the field.
As an overview, we summarize the takeaways of this paper below:
Overview of takeaways (Sec. 2.1) Template is crucial to make base models answer questions instead of completing sentences. In addition, all base models already possess math-solving capability prior to RL. (Sec. 2.2) Intriguingly, Qwen-2.5 base models get an immediate $\sim 60\%$ improvement by not using template, making us hypothesize that they may pretrain on concatenated question-answer texts when cooking the models. (Sec. 2.3) Nearly all base models already exhibit the “Aha moment”, including DeepSeek-V3-Base. (Sec. 3.1, Sec. 3.2) Dr. GRPO effectively fixes GRPO’s bias in optimization, achieving better token efficiency. (Sec. 3.3) Model-template mismatch can destruct reasoning capabilities before RL reconstructing it. (Sec. 3.4) Math pretraining on Llama-3.2-3B improves its RL ceiling.
@@ 2.1 R1-Zero Trainability: Templates Construct Exploratory Base Policies
Figure 3: Model attributes across three aspects. Question-Answering Ability: the extent to which a pretrained language model provides a direct answer to a question rather than continuing or expanding upon it; Exploration Ability: pass@8 measures how well base models explore; Self-Reflection: counts are obtained through cross-validation between keyword-based detection and LLM-based detection, as detailed in Appendix C.
Results. The left plot of Fig. 3 shows how well base models (with or without templates) answer the provided questions. We observe that Llama and DeepSeek models all improve the answering ability by employing the proper template (R1 template). However, Qwen2.5 models work best (with $100\%$ answering rate) when no template is used. This intriguing property motivates further investigation, as discussed in Sec. 2.2. Meanwhile, the lowest answering rate with no template suggests that DeepSeek-V3-Base is a nearly pure base model. This observation motivates us to explore whether a pure base model like DeepSeek-V3-Base demonstrates the Aha moment (Sec. 2.3). The middle plot of Fig. 3 shows the pass@8 accuracy of different base models (with template) at different sampling temperatures. This metric can serve as an indicator of base policy’s exploration ability. For example, if a base policy cannot even sample a single trajectory that leads to the correct final answer, it is impossible for RL to improve the policy because there is no reward signal. Our results demonstrate that all tested models are exploratory (thus ready for RL), with Qwen2.5 models performing the best (even surpassing DeekSeek-V3-Base). This might partially explain that most R1-Zero projects (Zeng et al., 2025; Hu et al., 2025) are based on Qwen2.5 models.
@@
We next dig into the intriguing observation (c.f. Fig. 3(Left)) that all Qwen2.5 base models readily serve as chat models even without any template. We take a step further to evaluate the reasoning ability of Qwen2.5-Math models on five standard benchmarks: AIME 2024 (Li et al., 2024), AMC (Li et al., 2024), MATH500 (Hendrycks et al., 2021), Minerva Math (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024). Following common practice, we use greedy decoding and limit the sampling budget to 3000 tokens.
@@ 2.3 Aha Moment Already Appears in Base Models Including DeepSeek-V3-Base
One of the most inspiring results of DeepSeek-R1-Zero is the emergence of self-reflection behaviors, a.k.a., Aha moment, through pure RL training. A few prior studies (Liu et al., 2025b; Yeo et al., 2025) have suggested that there may not be Aha moment in open-source R1 replications because the base models they use already exhibit self-reflection keywords. However, they have not tested DeepSeek-V3-Base, on which the real R1-Zero model was RL-tuned. We complete this missing piece by hosting DeepSeek-V3-Base-685B ourselves and investigating its responses of $500$ MATH questions with the R1 template. From the right plot of Fig. 3, we can observe that DeepSeek-V3-Base also generates a decent amount of self-reflections, further validating claims by Liu et al. (2025b). We also show examples in Fig. 4 where DeepSeek-V3-Base generates keywords such as “Aha”, “wait”, and “verify the problem”.
An additional important question is whether self-reflection behaviors are associated with improved model performance after RL training. To investigate this, we host DeepSeek-R1-Zero and analyze its responses of the same questions from MATH dataset. While self-reflection behaviors occur more frequently in R1-Zero, we observe these behaviors are not necessarily imply higher accuracy. Detailed analysis can be found in App. D.
@@ 3.1 GRPO Leads to Biased Optimization
Figure 5: Illustration of the biases in GRPO. Note that the effective advantage of GRPO $a_{i,t}$ is equivalent to a reweighted version of the unbiased advantage $\tilde{A}_{i,t}=R({\mathbf{q}},{\mathbf{o}}_{i})-\operatorname{mean}(\mathbf{R})$. The terms $\operatorname{std}(\mathbf{R})$ and $|{\mathbf{o}}_{i}|$ could bias the optimization by assigning different weights to different questions and responses, as denoted by the sizes of the blue circles and the lengths of the orange arrows.
@@
$R({\mathbf{q}},{\mathbf{o}})=\begin{cases}1&\text{if ${\mathbf{o}}$ contains the correct final answer to ${\mathbf{q}}$}\\ 0&\text{otherwise}\end{cases}$
We run RL on questions sampled from the MATH (Hendrycks et al., 2021) training dataset, and compare the vanilla GRPO with the proposed Dr. GRPO. We evaluate the online model on five benchmarks: AIME2024, AMC, MATH500, Minerva Math and OlympiadBench. More experimental details including hyperparameters can be found in our open-sourced codebase.
@@ 3.3 A Duet of Template and Question Set Coverage in RL dynamics
Question set Description
ORZ Combining AIME, Numina-Math, Tulu3 MATH; diverse and large amount (57k)
MATH High-school math competition questions (12k)
GSM Simpler grade-school math questions (8k)
[…]
@@
Results. Fig. 7 shows the RL curves of different runs, from which we can make several interesting observations: 1) Templates determine the performance of the initial policies, but RL can improve all policies to a comparable performance of ~$40\%$ (given a proper question set); 2) When using the R1 template, question sets have a significant impact on the dynamics of RL, with too narrow coverage leading to lower plateau performance. However, when using the Qwen-Math template, the best final performance is attained by RL on GSM-8K, demonstrating that training on much simpler (and o.o.d.) questions can largely improve (nearly double) the test accuracy on harder questions. From these observations, we draw the following insights:
  • The Qwen2.5-Math-1.5B base model already possesses strong math-solving capabilities (see the starting point in the right plot of Fig. 7). Applying templates in fact destructs the capability before RL reconstructing it. This implies that we shall be more conservative in claiming the huge gains brought about by pure RL.
  • When there is a large mismatch between base models and templates (e.g., R1 template mismatches Qwen2.5-Math-1.5B), the policy improvement mainly comes from RL-tuning, thus requiring question set to have good coverage (left plot of Fig. 7). Otherwise, even a small and completely o.o.d. question set could induce the reasoning ability equally well, by reinforcing correct reasoning behaviors instead of infusing new knowledge.
@@ 3.4 Domain-Specific Pretraining Improves RL Ceiling
Experimental settings. We adopt the Llama-3.2-3B base model as our starting point, and use the unbiased Dr. GRPO algorithm for RL-tuning with the R1 template. We hypothesize that domain-specific pretraining would help RL, hence we adopt the Llama-3.2-3B-FineMath [^4] https://huggingface.co/HuggingFaceTB/FineMath-Llama-3B., which is continual pretrained on the FineMath dataset (Allal et al., 2025). Moreover, as we hypothesize that Qwen2.5 models are likely to be pretrained on concatenated question-response texts (Sec. 2.2), we similarly prepare a concatenated dataset from NuminaMath-1.5 (LI et al., 2024), and continual pretrain Llama-3.2-3B-FineMath for 2 epochs with learning rate 1e-5. We refer to the concatanated continual pretrained model as Llama-3.2-3B-NuminaQA.
@@ Appendix A Policy Gradient Derivations
[…]
@@
[…]
@@ Appendix B Detailed Benchmark Results
Table 4: A comparison on benchmark scores. Oat-Zero models are RL-tuned by our minimalist recipe (Sec. 1). * means we employ the best template (no template) to generate answers, such that the test scores are highest and can faithfully reflect the capabilities of the base models.
Appendix C Keyword-based Detection and LLM-Based Identification of Self-Reflection Behaviors
@@
Appendix D Comparison Between DeepSeek-V3-Base and DeepSeek-R1-Zero
@@
Appendix E Prompts Used for GPT-As-A-Judge
@@
LLM-based Detection for Self-Reflection I will send you a mathematical question along with a detailed response. Your task is to determine whether the response is attempting to answer the question. If the response is off-topic, hallucinated, random talk, or otherwise irrelevant, mark it as 0. Otherwise, assess whether the response exhibits self-reflection. Categorization Rules: 1. Category 0: The response is off-topic, nonsensical, incoherent, overly repetitive, or lacks logical reasoning. Example cases: The response does not relate to the question. It contains meaningless or hallucinated content. It consists of excessive repetition without coherence. 2. Category 1: The response attempts to answer the question but does not exhibit self-reflection. Example cases: The response directly solves the problem without revisiting steps. No attempt is made to verify the correctness of the answer or explore alternative solutions. 3. Category 2: The response demonstrates self-reflection at any level. This may include: Explicit self-reflection keywords, such as: *recheck, rethink, reassess, reevaluate, re-evaluate, reevaluation, re-examine, reexamine, reconsider, reanalyze, double-check, check again, think again, verify again, go over the steps*, etc. Implicit self-reflection behaviors, such as revisiting the solution, questioning assumptions, or considering alternative approaches without explicit keywords. If any form of self-reflection is present, always categorize it as 2, regardless of correctness or answer quality. 4. Category 3: The response consists solely of Python code for calculations without exhibiting self-reflection. Example cases: The response only provides a Python script to compute the solution without any verification, re-evaluation, or alternative considerations. Output Format: Your response should first provide a very brief explanation of your analysis, followed by a single category number (0, 1, 2, or 3) at the end. You must include the category number at the end of your response. Example outputs: ‘The response is off-topic and does not attempt to answer the question. 0.‘ ‘The response provides a direct solution without self-reflection. 1.‘ ‘The response demonstrates self-reflection. 2.‘ ‘The response consists solely of Python code without any self-reflection. 3.‘ Question: {question} Response: {response}
Prompt for checking the model’s question-answering ability.
Prompt for Checking Question-Answering Ability I will send you a question and a long response generated by an LLM. Your task is to determine whether the output attempts to answer the question or not. The output may sometimes include irrelevant content, hallucinations, or random, off-topic responses. Please classify the output into one of the following categories: Output Format: Your response must start with a single integer (0 or 1), followed by a brief explanation. Return 0: → The output is not trying to answer the question (e.g., irrelevant content, random talking, hallucinations). Example output: ‘0: The response is off-topic and does not address the question.‘ Return 1: → The output attempts to answer the question, regardless of how complete or accurate the answer is. Example output: ‘1: The response engages with the question, even if the answer is incomplete or incorrect.‘ Question: {question} Response: {response}

### v2

@@ 1 Introduction
Our analysis on base models and RL suggests a minimalist recipe for R1-Zero-like training: we RL-tune Qwen2.5-Math-7B using the (unbiased) Dr. GRPO algorithm on MATH (Hendrycks et al., 2021) level 3-5 questions with the Qwen-Math template, and achieve state-of-the-art performance (Fig. 2) with only $27$ hours compute on $8\times$ A100 GPUs. We hope our findings presented in this paper, models released, and the codebase open-sourced could benefit future research in the field. As an overview, we summarize the takeaways of this paper below:
Overview of takeaways (Sec. 2.1) Template is crucial to make base models answer questions instead of completing sentences. In addition, all base models already possess math-solving capability prior to RL. (Sec. 2.2) Intriguingly, Qwen-2.5 base models get an immediate $\sim 60\%$ improvement by not using template, making us hypothesize that they may pretrain on concatenated question-answer texts when cooking the models. (Sec. 2.3) Nearly all base models already exhibit the “Aha moment”, including DeepSeek-V3-Base. (Sec. 3.1, Sec. 3.2) Dr. GRPO effectively fixes GRPO’s bias in optimization, achieving better token efficiency. (Sec. 3.3) Model-template mismatch can destroy reasoning capabilities before RL reconstructs it. (Sec. 3.4) Math pretraining on Llama-3.2-3B improves its RL ceiling.
@@ 2.1 R1-Zero Trainability: Templates Construct Exploratory Base Policies
Figure 3: Model attributes across three aspects. Question-Answering Ability: the extent to which a pretrained language model provides a direct answer to a question rather than continuing or expanding upon it; Exploration Ability: pass@8 measures how well base models explore; Self-Reflection: counts are obtained through cross-validation between keyword-based detection and LLM-based detection, as detailed in Appendix D.
Results. The left plot of Fig. 3 shows how well base models (with or without templates) answer the provided questions. We observe that Llama and DeepSeek models all improve the answering ability by employing the proper template (R1 template). However, Qwen2.5 models work best (with $100\%$ answering rate) when no template is used. This intriguing property motivates further investigation which will be discussed in Sec. 2.2. Meanwhile, the lowest answering rate with no template suggests that DeepSeek-V3-Base is a nearly pure base model. This observation motivates us to explore whether a pure base model like DeepSeek-V3-Base demonstrates the Aha moment (Sec. 2.3). The middle plot of Fig. 3 shows the pass@8 accuracy of different base models (with template) at different sampling temperatures. This metric can serve as an indicator of base policy’s exploration ability. For example, if a base policy cannot even sample a single trajectory that leads to the correct final answer, it is impossible for RL to improve the policy because there is no reward signal. Our results demonstrate that all tested models are exploratory (thus ready for RL), with Qwen2.5 models performing the best (even surpassing DeekSeek-V3-Base). This might partially explain that most R1-Zero projects (Zeng et al., 2025; Hu et al., 2025) are based on Qwen2.5 models.
@@
We next dig into the intriguing observation (c.f. Fig. 3(Left)) that all Qwen2.5 base models readily serve as chat models even without any template. We take a step further to evaluate the reasoning ability of Qwen2.5-Math models on five standard benchmarks: AIME 2024 (Li et al., 2024a), AMC (Li et al., 2024a), MATH500 (Hendrycks et al., 2021), Minerva Math (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024). Following common practice, we use greedy decoding and limit the sampling budget to 3000 tokens.
@@ 2.3 Aha Moment Already Appears in Base Models Including DeepSeek-V3-Base
One of the most inspiring results of DeepSeek-R1-Zero is the emergence of self-reflection behaviors, a.k.a., Aha moment, through pure RL training. A few prior studies (Liu et al., 2025b; Yeo et al., 2025) have suggested that there may not be Aha moment in open-source R1 replications because the base models they use already exhibit self-reflection keywords. However, they have not tested DeepSeek-V3-Base, on which the real R1-Zero model was RL-tuned. We complete this missing piece by hosting DeepSeek-V3-Base-685B ourselves and investigating its responses to the $500$ MATH questions with the R1 template. From the right plot of Fig. 3, we can observe that DeepSeek-V3-Base also generates a decent amount of self-reflections, further validating the claims of Liu et al. (2025b). We also show examples in App. E (Fig. 13) where DeepSeek-V3-Base generates keywords such as “Aha” and “wait”.
An additional important question is whether self-reflection behaviors are associated with improved model performance after RL training. To investigate this, we host DeepSeek-R1-Zero and analyze its responses to the same questions from the MATH dataset. Although self-reflection behaviors occur more frequently in R1-Zero, we observe that these behaviors are not positively correlated with higher accuracy. Detailed analysis can be found in App. F.
@@ 3.1 GRPO Leads to Biased Optimization
Figure 4: Illustration of the biases in GRPO. Note that the effective advantage of GRPO $a_{i,t}$ is equivalent to a reweighted version of the unbiased advantage $\tilde{A}_{i,t}=R({\mathbf{q}},{\mathbf{o}}_{i})-\operatorname{mean}(\mathbf{R})$. The terms $\operatorname{std}(\mathbf{R})$ and $|{\mathbf{o}}_{i}|$ could bias the optimization by assigning different weights to different questions and responses, as denoted by the sizes of the blue circles and the lengths of the orange arrows. Upward arrows indicate positive advantages, and vice versa.
@@
[…]
We run RL on questions sampled from the MATH (Hendrycks et al., 2021) training dataset, and compare the vanilla GRPO with the proposed Dr. GRPO. We evaluate the online model on five benchmarks: AIME2024, AMC, MATH500, Minerva Math and OlympiadBench. More experimental details including hyperparameters can be found in App. G.
@@ 3.3 A Duet of Template and Question Set Coverage in RL dynamics
Question set # Description
ORZ 57k Combining AIME, Numina-Math, Tulu3 MATH; diverse and large amount
MATH 12k High-school math competition questions
GSM 8k Simpler grade-school math questions
ASDiv 2k Basic algebra ($+-\times\div)$ questions
@@
Results. Fig. 6 shows the RL curves of different runs, from which we can make several interesting observations: 1) Templates determine the performance of the initial policies, but RL can improve all policies to a comparable performance of $\sim 40\%$ (given a proper question set); 2) When using the R1 template, question sets have a significant impact on the dynamics of RL, with too narrow coverage leading to lower plateau performance. However, when using the Qwen-Math template, the best final performance is attained by RL on GSM-8K, demonstrating that training on much simpler (and o.o.d.) questions can largely improve (nearly double) the test accuracy on harder questions. From these observations, we draw the following insights:
  • The Qwen2.5-Math-1.5B base model already possesses strong math-solving capabilities (see the starting point in the right plot of Fig. 6). Applying templates in fact destroys the capability before RL reconstructs it. This implies that we should be more conservative in claiming the huge gains brought about by pure RL.
  • When there is a large mismatch between base models and templates (e.g., R1 template mismatches Qwen2.5-Math-1.5B), the policy improvement mainly comes from RL-tuning, thus requiring question set to have good coverage (left plot of Fig. 6). Otherwise, even a small and completely o.o.d. question set could induce the reasoning ability equally well, by reinforcing useful reasoning behaviors instead of infusing new knowledge.
@@ 3.4 Domain-Specific Pretraining Improves RL Ceiling
Experimental settings. We adopt the Llama-3.2-3B base model as our starting point, and use the unbiased Dr. GRPO algorithm for RL-tuning with the R1 template. We hypothesize that domain-specific pretraining would help RL, hence we adopt the Llama-3.2-3B-FineMath [^4] https://huggingface.co/HuggingFaceTB/FineMath-Llama-3B., which is continual pretrained on the FineMath dataset (Allal et al., 2025). Moreover, as we hypothesize that Qwen2.5 models are likely to be pretrained on concatenated question-response texts (Sec. 2.2), we similarly prepare a concatenated dataset from NuminaMath-1.5 (Li et al., 2024b), and continual pretrain Llama-3.2-3B-FineMath for 2 epochs with learning rate 1e-5. We refer to the concatanated continual pretrained model as Llama-3.2-3B-NuminaQA.
@@ Appendix A Policy Gradient Derivations
$\tilde{A}_{i,t}=\frac{R({\mathbf{q}},{\mathbf{o}}_{i})-\operatorname{mean}({\{R({\mathbf{q}},{\mathbf{o}}_{1}),\dots,R({\mathbf{q}},{\mathbf{o}}_{G})\}})}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\xcancel{\operatorname{std}({\{R({\mathbf{q}},{\mathbf{o}}_{1}),\dots,R({\mathbf{q}},{\mathbf{o}}_{G})\}})}}}.$
@@
$\begin{split}\mathcal{J}(\pi_{\theta})&=\mathbb{E}{[{\mathbf{q}}\sim p_{\mathcal{Q}},\{{\mathbf{o}}_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|{\mathbf{q}})]}\\ &\frac{1}{G}\sum_{i=1}^{G}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\xcancel{\frac{1}{|{\mathbf{o}}_{i}|}}}\sum_{t=1}^{|{\mathbf{o}}_{i}|}\left\{\min\left[\frac{\pi_{\theta}(o_{i,t}|{\mathbf{q}},{\mathbf{o}}_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}|{\mathbf{q}},{\mathbf{o}}_{i,<t})}\tilde{A}_{i},\text{clip}\left(\frac{\pi_{\theta}(o_{i,t}|{\mathbf{q}},{\mathbf{o}}_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}|{\mathbf{q}},{\mathbf{o}}_{i,<t})},1-\epsilon,1+\epsilon\right)\tilde{A}_{i}\right]\right\},\end{split}$
@@ Appendix B Detailed Benchmark Results
Table 4: A comparison on benchmark scores. Ours models are RL-tuned by our minimalist recipe (Sec. 1). * means we employ the best template (no template) to generate answers, such that the test scores are highest and can faithfully reflect the capabilities of the base models.
Appendix C Extended Empirical Results
In this section we present two extended empirical results for (1) the ablation of different bias terms in GRPO and (2) statistical significance of Dr. GRPO’s results. We RL-tune the Qwen2.5-1.5B base model on a mixture of 3K diverse math questions drawn from ASDiv, MATH, and AIME (pre-2023).
Figure 8: Ablation results on the two bias terms in GRPO.
Fig. 8 shows the training and evaluation curves for the following variants: Dr. GRPO, GRPO w/o length normalization, GRPO w/o standard deviation (std) normalization and Vanilla GRPO. From the middle subplot, we observe that both Dr. GRPO and the variant without length normalization generate shorter responses compared to the other two. This confirms that the length bias term has a more significant influence on response length–consistent with our expectations.
In terms of performance, Dr. GRPO and the other ablated variants consistently outperform vanilla GRPO in both training rewards and evaluation accuracy. This indicates that removing bias terms (either length or std) improves policy learning, validating our motivation for Dr. GRPO.
Figure 9: Evaluation results of 3 independent RL runs. The mean curves are drawn in solid lines and the standard deviation is plotted in the shaded areas.
Fig. 9 compares GRPO and Dr. GRPO across three independent runs. We observe that Dr. GRPO consistently demonstrates statistically significant improvements–both in token efficiency and final accuracy–across different random seeds.
Appendix D Keyword-based Detection and LLM-Based Identification of Self-Reflection Behaviors
@@
Appendix E Examples of Aha Moment in DeepSeek-V3-Base
Fig. 13 shows two examples to demonstrate that the DeepSeek-V3-Base model already exhibits the so-called “aha moment” even before the RL-tuning.
Appendix F Comparison Between DeepSeek-V3-Base and DeepSeek-R1-Zero
@@
Appendix G Detailed Experimental Settings
All our experiments are performed on 8 $\times$ A100 GPUs and finished in about one day. We enable the actor-learner collocation supported by Oat (Liu et al., 2025a) to optimize the training efficiency. We show the experimental configurations in Table 6.
Parameter Value
Actor
Maximum response length $3000$ tokens
Sampling temperature 1.0
(top P, top k) (1.0, -1)
Number of responses per question 8
Learner
Optimizer AdamW
Adam parameters ($\beta_{1},\beta_{2}$) (0.9, 0.95)
Weight decay 0.0
Gradient norm clipping 1.0
Learning rate scheduler Constant
Learning rate $1\times 10^{-6}$
Inner proximal update epoch 1
KL loss coefficient 0.0
KL penalty coefficient 0.0
Policy clipping parameter 0.2
Table 6: Hyperparameter configurations used in all experiments.
Appendix H Prompts Used for GPT-As-A-Judge
Prompt for checking the model’s question-answering ability.
Prompt for Checking Question-Answering Ability I will send you a question and a long response generated by an LLM. Your task is to determine whether the output attempts to answer the question or not. The output may sometimes include irrelevant content, hallucinations, or random, off-topic responses. Please classify the output into one of the following categories: Output Format: Your response must start with a single integer (0 or 1), followed by a brief explanation. Return 0: → The output is not trying to answer the question (e.g., irrelevant content, random talking, hallucinations). Example output: ‘0: The response is off-topic and does not address the question.‘ Return 1: → The output attempts to answer the question, regardless of how complete or accurate the answer is. Example output: ‘1: The response engages with the question, even if the answer is incomplete or incorrect.‘ Question: {question}
Response: {response}
@@ Appendix H Prompts Used for GPT-As-A-Judge
LLM-based Detection for Self-Reflection I will send you a mathematical question along with a detailed response. Your task is to determine whether the response is attempting to answer the question. If the response is off-topic, hallucinated, random talk, or otherwise irrelevant, mark it as 0. Otherwise, assess whether the response exhibits self-reflection. Categorization Rules: 1. Category 0: The response is off-topic, nonsensical, incoherent, overly repetitive, or lacks logical reasoning. Example cases: The response does not relate to the question. It contains meaningless or hallucinated content. It consists of excessive repetition without coherence. 2. Category 1: The response attempts to answer the question but does not exhibit self-reflection. Example cases: The response directly solves the problem without revisiting steps. No attempt is made to verify the correctness of the answer or explore alternative solutions. 3. Category 2: The response demonstrates self-reflection at any level. This may include: Explicit self-reflection keywords, such as: *recheck, rethink, reassess, reevaluate, re-evaluate, reevaluation, re-examine, reexamine, reconsider, reanalyze, double-check, check again, think again, verify again, go over the steps*, etc. Implicit self-reflection behaviors, such as revisiting the solution, questioning assumptions, or considering alternative approaches without explicit keywords. If any form of self-reflection is present, always categorize it as 2, regardless of correctness or answer quality. 4. Category 3: The response consists solely of Python code for calculations without exhibiting self-reflection. Example cases: The response only provides a Python script to compute the solution without any verification, re-evaluation, or alternative considerations. Output Format: Your response should first provide a very brief explanation of your analysis, followed by a single category number (0, 1, 2, or 3) at the end. You must include the category number at the end of your response. Example outputs: ‘The response is off-topic and does not attempt to answer the question. 0.‘ ‘The response provides a direct solution without self-reflection. 1.‘ ‘The response demonstrates self-reflection. 2.‘ ‘The response consists solely of Python code without any self-reflection. 3.‘ Question: {question}
Response: {response}